# RISK: A Framework for GUI Agents in E-commerce Risk Management

### v1

@@ 4.2 Reward Design
In general GUI agents, the reward function typically focuses on the grounding accuracy of actions (e.g., the predicted point should be within the bounding box of the target element). However, this kind of reward function cannot satisfy the requirements of e-commerce risk management tasks due to complex webpages and diverse task scenarios. As shown in Figure 4, when there are dense elements on the page, the DOM tree structure is more suitable for accurate interaction. Therefore, a more dependable and comprehensive reward function is needed to guide the learning process of GUI agents in this domain. We detail our design below.
@@
Format Reward. Format reward $R_{for}$ is introduced to ensure the syntactic and semantic correctness of the model’s output. As ‘think’ content and ‘action’ content are still required in the output, we also consider the ‘evaluation_previous_goal’, ‘memory’, and ‘next_goal’ content, which come from the Browser Use framework and are beneficial for the model to understand the task process and webpage status. Among them, ‘evaluation_previous_goal’ is used to evaluate whether the last step’s action is completed, ‘memory’ records the current task status, and ‘next_goal’ describes the next step’s action. Moreover, since that RISK-R1 does not employ (x,y) coordinates for element selection but uses the element index in the DOM tree combined with the tools, we additionally check the correctness of the ‘action’ content format, which should be in the form of ‘[{<tool_name>:{<index>,<text>(optional)}}]’. This design ensures that the model’s output is well-structured and interpretable, facilitating practical application in real scenarios. The format reward $R_{for}$ is 1 if the output format is correct, and 0 otherwise.
@@
Stepwise Accuracy Reward. In practical tool calling scenarios, the tool list in the action predicted by the model may contain multiple actions to save the number of MLLM calls. Considering that the nature of RFT is to assist the model in exploring the correct path, the original binary accuracy reward treating the entire tool list as a whole is too coarse-grained to provide effective guidance at the early stage of training. Therefore, we propose a stepwise accuracy reward $R_{step\_acc}$ that evaluates the accuracy of each action in the list, providing more detailed feedback to the model at the early stage of training to facilitate exploration. After training the model to a certain extent, we further fine-tune it with the original binary accuracy reward to avoid the model exhibiting inertia under partial rewards. Specifically, for a tool list $T=\{t_{1},t_{2},\cdots,t_{n}\}$ in the action, where $t_{i}$ is the $i$-th tool in the list, we define the stepwise accuracy reward $R_{step\_acc}$ as follows:
$R_{step\_acc}=\begin{cases}\frac{1}{n}\sum_{i=1}^{n}R_{acc}(t_{i})&\text{early stage},\\ R_{acc}(T)&\text{later stage},\\ \end{cases}\qquad R_{acc}(t_{i})=\begin{cases}1&\text{if }F_{1}(t_{i},t_{i}^{gt})>0.5,\\ 0&\text{otherwise},\end{cases}$ (2)
where $F_{1}(t_{i},t_{i}^{gt})$ is the F1 score between the predicted tool $t_{i}$ and the ground truth tool $t_{i}^{gt}$.
@@
Process Reweight. The motivation for process reweight comes from two aspects: (1) In business-oriented GUI tasks, the initial steps and associated webpage content are relatively simple (e.g., opening the Google search page), whereas later steps involve more complex pages (e.g., specific e-commerce pages), and (2) Early-stage steps exhibit high homogeneity, while later-stage steps show greater differentiation. Therefore, we propose a process reweighting $\theta$ to distinguish the importance of different steps in a trajectory, emphasizing the later steps that are more critical for task completion. We design a weight curve that increases with the step index, where the weight of the first step is $\gamma$ and the last step is 1.0, as shown in Appendix Figure 9. The process reweight $\theta$ for the $i$-th step in a trajectory with $n$ steps is defined as follows: $\theta(i)=\gamma+(1-\gamma)\left(1+e^{-\left(2\delta\frac{i-1}{n-1}-\delta\right)}\right)^{-1},$ where $\gamma$ and $\delta$ are hyperparameters to control the shape of the curve.
@@ 4.3 Reinforcement Learning Objective
To leverage the advantages of each component in the reward design, we combine them to form the overall reward $R$ for RISK-R1:
$R=\alpha\cdot R_{for}+\beta\cdot\theta\cdot R_{step\_acc}$ (3)
where $\alpha$ and $\beta$ are hyperparameters to balance the contributions of each component. Based on the overall reward $R$, we compute the advantage $A$ and optimize the policy model using the GRPO objective in Equation 1.
@@
Level Reweight. In RISK-Data, the samples are graded into three difficulty levels: easy, moderate, and difficult. As demonstrated in (Zhou et al., 2025), question-level difficulty bias is beneficial for the model to focus on challenging aspects of the task. Therefore, we introduce a level reweight $w_{level}$ to adjust the contribution of samples at different difficulty levels to the objective function. Specifically, we set the level reweight $w_{level}$ for easy, moderate, and difficult samples as 1.0, 1.1, and 1.2, respectively. The final optimization objective of RISK-R1 is modified to $w_{level}\cdot\mathcal{J}_{GRPO}(\pi_{\theta})$. Previously, the relative grounding box size was used to set the difficulty level, which is not appropriate since this criterion only considers the element size in isolation but ignores the element density and page complexity. This limitation may lead to a wrong assessment of task difficulty, where empirical evidence is shown in Table 3. In contrast, our difficulty grading based on the advanced MLLM’s accuracy is more comprehensive and dependable, where the comparison of KL divergence curves under different level reweight settings at early training stages is shown in Figure 5. It can be observed that the model with level reweighting deviates more and faster from the reference model, indicating that it explores more diverse strategies and learns more effectively from the challenging samples.
@@ 5.1 Experimental Setup
Implementation Details. For SFT, we use the Qwen2.5-VL-7B-Instruct as the base model and train it for one epoch to learn the basic interaction capabilities. For RFT, we initialize the policy model with the supervised fine-tuned model and use the VeRL framework (Sheng et al., 2024) for training over six epochs. RFT Training is conducted on 8 NVIDIA H200-141G GPUs with the following hyperparameters: learning rate of 1e-6, rollouts per prompt of 8, and KL coefficient of 0.04. As the format has been initially standardized in SFT, we set reward coefficients $\alpha=0.1$ and $\beta=0.9$. The default process reweight coefficients are set to $\gamma=0.7$ and $\delta=4$. We use a stepwise reward in the first epoch and a binary reward in the remaining epochs.
Training Datasets and Evaluation Benchmarks. In SFT, we use all single-step and multi-step trajectories in RISK-Data for training. In RFT, we only use the single-step trajectories since the multi-step trajectories are too long to fit in the GPU memory. Considering general grounding data is beneficial for improving the model’s website perception and element manipulation capabilities, we also incorporate the GUI-R1 (Luo et al., 2025) dataset into our training data. We evaluate RISK-R1 from three aspects: (1) Offline evaluation on RISK-Bench to assess the model’s performance in e-commerce risk management tasks, (2) Offline evaluation on general GUI navigation benchmark OS-Genesis (Sun et al., 2024) to evaluate the model’s generalization ability, where the web tasks are tested, and (3) Online evaluation in real-world e-commerce risk management scenarios to validate the practical effectiveness of RISK-R1.
We elaborate more experimental details in Appendix A.6.
@@ 5.2 Main Results
Table 1: Performance comparison of different models on RISK-Bench and OS-Genesis. The best results and the second best results are highlighted in bold and underline, respectively. Our RISK-R1-7B achieves SOTA performance, surpassing all baselines across single-step and multi-step tasks.
Model RISK-Bench OS-Genesis Single-step Multi-step $\uparrow$ Web Task $\uparrow$ Easy $\uparrow$ Moderate $\uparrow$ Difficult $\uparrow$ Overall $\uparrow$ Commercial Models GPT-4o 98.2 82.9 46.8 81.5 74.0 55.3 Qwen-VL-Max 95.8 78.5 22.4 72.9 50.0 50.3 General Open-source Models InternVL3-8B 30.1 14.3 0.0 18.8 0.0 29.5 Qwen2.5-VL-7B 62.3 45.4 4.8 43.6 0.6 32.2 Qwen2.5-VL-72B 98.8 81.9 42.9 80.6 67.8 50.0 GUI-specific Models (SFT) UI-TARS-2B 0.2 0.0 0.0 0.1 0.0 1.3 UI-TARS-7B 11.3 5.5 0.0 7.1 0.0 4.2 UI-TARS-72B 20.7 9.3 0.9 13.0 0.0 5.8 OS-Atlas-7B 37.3 27.7 2.4 26.1 0.0 23.0 ShowUI-2B 6.9 2.7 0.0 4.2 0.0 3.4 Aguvis-7B 8.3 4.5 0.0 5.3 0.0 29.7 GUI-specific Models (RL) GUI-R1-3B 55.4 37.2 3.0 37.8 0.0 24.3 GUI-R1-7B 65.4 45.0 9.3 46.3 0.0 28.0 InfiGUI-R1-3B 18.6 11.7 1.4 12.6 0.0 10.6 GUI-G1-3B 55.1 38.9 4.5 38.4 0.0 19.1 UI-TARS-1.5-7B 44.9 28.4 1.9 30.1 0.0 26.5 UI-Venus-Navi-7B 0.7 0.0 0.0 0.3 0.0 14.3 Ours RISK-SFT-7B 99.1 83.2 52.5 83.5 75.3 61.5 RISK-R1-7B 98.8 90.1 65.5 88.3 82.8 62.3
Table 2: Difficulty measurement analysis.
Measurement Single-step Multi-step
No Reweighting 86.7 79.6
Rule Score 86.1 (-0.6) 78.0 (-1.6)
LLM Response 88.3 (+1.6) 82.8 (+4.8)
Table 3: Difficulty weight configurations.
Configuration Single-step Multi-step
{0.8,0.9,1.0} 87.8 82.0
{1.0,1.1,1.2} 88.3 82.8
{1.0,1.3,1.5} 88.1 82.4
Offline Domain Evaluation. We compare RISK-R1 with commercial models, general open-source models, and GUI-specific models on RISK-Bench and OS-Genesis, as shown in Table 1. RISK-R1-7B surpasses all baselines across single-step and multi-step tasks. Specifically, in single-step tasks on RISK-Bench, RISK-R1-7B attains an overall accuracy of 88.3%, outperforming GPT-4o by 6.8% and Qwen2.5-VL-72B by 7.7%. Notably, after RFT, RISK-R1-7B has a slight decrease of 0.3% in easy tasks but a substantial increase of 6.9% and 22.6% in moderate and difficult tasks, respectively. This change reveals that level reweighting effectively guides the model to focus on challenging samples, enhancing its problem-solving capabilities. In multi-step tasks, RISK-R1-7B achieves a task success rate of 82.8%, exceeding GPT-4o by 8.8%, indicating that multi-step trajectories in RISK-Data are beneficial for improving the model’s task-level process understanding, while process reweighting emphasizes the importance of later steps in the trajectory, further enhancing performance.
@@
Table 4: Performance comparison of models on RISK-Bench with online evaluation, where webpage interaction is built on Browser-Use. Compared with SOTA baselines, RISK-R1-7B achieves the highest task success rate while maintaining a competitive task completion rate.
Model Completion Success
Commercial Models
Qwen-VL-Max 85.2 66.2
General Open-source Models
InternVL3-8B 0.0 0.0
Qwen2.5-VL-7B 8.3 46.4
Qwen2.5-VL-72B 88.7 68.9
GUI-specific Models (SFT)
UI-TARS-7B 0.0 0.0
UI-TARS-72B 0.0 0.0
OS-Atlas-7B 4.4 37.2
ShowUI-2B 0.0 0.0
GUI-specific Models (RL)
GUI-R1-7B 0.0 0.0
InfiGUI-R1-3B 0.0 0.0
GUI-G1-3B 0.0 0.0
UI-TARS-1.5-7B 0.0 0.0
Ours
RISK-SFT-7B 86.1 67.0
RISK-R1-7B 87.6 70.5
Online Evaluation. For read-time multi-step decision-making evaluation, we use the Browser-Use framework to build a webpage interaction environment and compare RISK-R1 with various baselines. Different from offline evaluations, the model may encounter unseen webpages or changed page structures (even if the objective website is the same) during online evaluations, which poses a greater challenge to the model’s generalization ability and robustness. As shown in Table 4, although the task completion rate of RISK-R1-7B is slightly lower than that of Qwen2.5-VL-72B, it achieves the highest task success rate of 70.5%, outperforming Qwen2.5-VL-72B by 1.6% and Qwen-VL-Max by 4.3%. This indicates that RISK-R1-7B can effectively complete tasks even in complex and dynamic real-world scenarios, demonstrating its practical applicability in e-commerce risk management.
@@ 5.3 Reward Design Analysis
Level Reweight: Difficulty Grading Matters. In RISK-R1, we use the advanced MLLM’s accuracy to grade the difficulty of samples and set the level reweight accordingly. To validate the effectiveness of this grading method, we compare it with no reweighting and rule-based scoring methods, as shown in Table 3. The rule-based scoring method assigns difficulty levels based on the tool count at each step, where easy, moderate, and difficult samples contain 1, 2, and more than 2 tools, respectively. The results indicate that inappropriate difficulty grading methods can negatively impact model performance, with the rule-based scoring method leading to decreases of 0.6% and 1.6% in single-step and multi-step tasks, respectively. Another notable setting is the difficulty weight configuration, which influences the emphasis on challenging samples. As shown in Table 3, the configuration {1.0, 1.1, 1.2} yields the best performance, while both overly flat ({0.8, 0.9, 1.0}) and overly steep ({1.0, 1.3, 1.5}) configurations lead to performance drops. This analysis highlights the importance of appropriate difficulty grading and weight configuration in guiding the reinforcement learning optimization.
Figure 7: Stepwise accuracy reward analysis. Peak performance is achieved at the combination of stepwise and binary accuracy rewards.
Process Reweight: Critical Steps in Task Completion. The goal of process reweighting is to emphasize the importance of later steps in a trajectory, which are more distinguishable and determine the success of task completion. We conduct hyperparameter sensitivity analysis for process reweight, as shown in Figure 7. $\gamma$ controls the weight of the initial step and $\delta$ controls the growth rate of the weight curve, where the visualization is shown in Figure 9. Comparison results reveal that an appropriate setting of process reweight can guide the model to focus on critical steps. However, an excessively low $\gamma$ (e.g., 0.4) or an excessively high $\delta$ (e.g., 7) leads to performance degradation, as the model may neglect the importance of early and intermediate steps, leading to suboptimal learning outcomes.
@@
Stepwise Accuracy Reward: Fine-grained Feedback for Exploration. RISK-R1 employs stepwise accuracy reward in the early stage of training to provide more fine-grained feedback for the model, facilitating exploration. We analyze the impact of different reward settings, as shown in Figure 7 (All reward curves are smoothed). Leveraging stepwise accuracy reward in the early stage provides a faster reward enhancement by encouraging the model to learn from partially correct tool calls. Nevertheless, using stepwise accuracy reward throughout the entire training process does not yield better results than the solely binary accuracy reward, as the model may develop inertia under partial rewards. The optimal approach is to combine both reward types, using stepwise accuracy reward in the early stage and binary accuracy reward in the later stage, which effectively balances exploration and exploitation.

### v2

@@ 4.2 Reward Design
In general GUI agents, the reward function typically focuses on the grounding accuracy of actions (e.g., the predicted point should be within the bounding box of the target element). However, this cannot satisfy the requirements of e-commerce risk management tasks due to complex webpages and diverse task scenarios. As shown in Figure 3, when there are dense elements on the page, the DOM tree structure is more suitable for accurate interaction. Therefore, a more dependable and comprehensive reward function is needed to guide the learning process of GUI agents.
@@
Format Reward. Format reward $R_{for}$ is introduced to ensure the syntactic and semantic correctness of the model’s output. As ‘think’ content and ‘action’ content are still required in the output, we also consider the ‘evaluation_previous_goal’, ‘memory’, and ‘next_goal’ content, which come from the Browser Use framework and are beneficial for the model to understand the task process and webpage status. Among them, ‘evaluation_previous_goal’ is used to evaluate whether the last step’s action is completed, ‘memory’ records the current task status, and ‘next_goal’ describes the next step’s action. Moreover, since that RISK-R1 does not employ (x,y) coordinates for element selection but uses the element index in the DOM tree combined with the tools, we additionally check the correctness of the ‘action’ content format, which should be in the form of ‘[{<tool_name>:{<index>,<text>(optional)}}]’. This design ensures that the model’s output is well-structured and interpretable, facilitating practical applications. The format reward $R_{for}$ is 1 if the output format is correct, and 0 otherwise.
@@
Stepwise Accuracy Reward. In practical tool calling scenarios, the tool list in the action predicted by the model may contain multiple actions to save the number of MLLM calls. Considering that the nature of RFT is to assist the model in exploring the correct path, the original binary accuracy reward treating the entire tool list as a whole is too coarse-grained to provide effective guidance at the early stage of training. Therefore, we propose a stepwise accuracy reward $R_{step\_acc}$ that evaluates the accuracy of each action in the list, providing more detailed feedback to the model at the early stage of training to facilitate exploration. After training the model to a certain extent, we further fine-tune it with the original binary accuracy reward to avoid the model exhibiting inertia under partial rewards. Specifically, given an action tool list $T=\{t_{1},t_{2},\cdots,t_{n}\}$ where $t_{i}$ is the $i$-th tool, we define the early-stage stepwise accuracy reward as $R_{step\_acc}=\frac{1}{n}\sum_{i=1}^{n}R_{acc}(t_{i})$, and use the $R_{acc}(T)$ at the later stage, where $R_{acc}(t_{i})=\mathbb{I}[F_{1}(t_{i},t_{i}^{gt})>0.5]$ ($F_{1}$ is the F1 score). Detailed analysis is shown in Sec. 5.3.
@@
Process Reweight. The motivation for process reweight comes from two aspects: (1) In business-oriented GUI tasks, the initial steps and associated webpage content are relatively simple (e.g., opening the Google search page), whereas later steps involve more complex pages (e.g., specific e-commerce pages), and (2) Early-stage steps exhibit high homogeneity, while later-stage steps show greater differentiation. Therefore, we propose a process reweighting $\theta$ to distinguish the importance of different steps in a trajectory, emphasizing the later steps that are more critical for task completion. The process reweight $\theta$ for the $i$-th step in a trajectory with $n$ steps is defined as follows: $\theta(i)=\gamma+(1-\gamma)\left(1+e^{-\left(2\delta\frac{i-1}{n-1}-\delta\right)}\right)^{-1},$ where $\gamma$ and $\delta$ are hyperparameters. More experimental analysis are shown in Appx. G.1.
@@ 4.3 Reinforcement Learning Objective
To leverage the advantages of each component in the reward design, we combine them to form the overall reward $R$ for RISK-R1: $R=\alpha\cdot R_{for}+\beta\cdot\theta\cdot R_{step\_acc},$ where $\alpha$ and $\beta$ are hyperparameters to balance the contributions of each component. Based on the overall reward $R$, we compute the advantage $A$ and optimize the policy model using the GRPO objective in Equation 4.1.
@@
Level Reweight. In RISK-Data, the samples are graded into three difficulty levels: easy, moderate, and difficult. As demonstrated in (Zhou et al., 2025), question-level difficulty bias is beneficial for the model to focus on challenging aspects of the task. Therefore, we introduce a level reweight $w_{level}$ to adjust the contribution of samples at different difficulty levels to the objective function. The final optimization objective of RISK-R1 is modified to $w_{level}\cdot\mathcal{J}_{GRPO}(\pi_{\theta})$. Previously, the relative grounding box size was used to set the difficulty level, which is not appropriate since this criterion only considers the element size in isolation but ignores the element density and page complexity. This limitation may lead to a wrong assessment of task difficulty. In contrast, our difficulty grading based on the advanced MLLM’s accuracy is more comprehensive, where the comparison of KL divergence curves under different level reweight settings at early training stages is shown in Figure 4. The model with level reweighting deviates more and faster from the reference model, indicating that it explores more diverse strategies and learns more effectively from the challenging samples. More empirical analysis are shown in Appx. G.2.
@@ 5.1 Experimental Setup
Implementation Details. For SFT, we use the Qwen2.5-VL-7B-Instruct as the base model and train it for one epoch to learn the basic interaction capabilities. For RFT, we initialize the policy model with the SFT model and use the VeRL framework (Sheng et al., 2024) for training over six epochs. RFT Training is conducted on 8 NVIDIA H200-141G GPUs with the following hyperparameters: learning rate of 1e-6, rollouts per prompt of 8, and KL coefficient of 0.04. As the format has been initially standardized in SFT, we set reward coefficients $\alpha=0.1$ and $\beta=0.9$. The default process reweight coefficients are set to $\gamma=0.7$ and $\delta=4$. We use a stepwise reward in the first epoch and a binary reward in the remaining epochs.
Training and Evaluation. In SFT, we use all trajectories in RISK-Data for training. In RFT, we only use the single-step trajectories since the multi-step trajectories are too long to fit in the GPU memory. Considering general grounding data is beneficial for enhancing model’s website perception and element manipulation capabilities, we incorporate the GUI-R1 (Luo et al., 2025) dataset into our training data. We evaluate RISK-R1 from three aspects: (1) Offline evaluation on RISK-Bench to assess the model’s performance in e-commerce risk management tasks, (2) Offline evaluation on general GUI navigation benchmark OS-Genesis (Sun et al., 2024) to evaluate the model’s generalization ability, where the web tasks are tested, and (3) Online evaluation in real-world e-commerce risk management scenarios to validate the practical effectiveness of RISK-R1. More details are provided in Appx. F.
Evaluation Metrics. In offline single-step trajectory evaluations, we use the accuracy of tool calls as the evaluation metric, where a tool call is considered correct if its F1 score with the ground truth tool call exceeds 0.5. In offline multi-step trajectory evaluations, we use the task success rate as the evaluation metric, where a trajectory is considered successfully completed if all tool calls in the trajectory are correct. In online evaluations, we use the task completion rate and task success rate as evaluation metrics, where the task completion rate is the percentage of tasks completed within a limited step counts (set to 20), and the task success rate is the percentage of tasks answered successfully.
@@ 5.2 Main Results
Table 1: Performance comparison (success rate) of different models on RISK-Bench and OS-Genesis. The best and the second best results are highlighted in bold and underline, respectively. Our RISK-R1-7B surpasses all baselines across single-step and multi-step tasks.
Model RISK-Bench OS-Genesis Single-step Multi-step Web Task Easy Moderate Difficult Overall Commercial Models GPT-4o 98.2 82.9 46.8 81.5 74.0 55.3 Qwen-VL-Max 95.8 78.5 22.4 72.9 50.0 50.3 General Open-source Models InternVL3-8B 30.1 14.3 0.0 18.8 0.0 29.5 Qwen2.5-VL-7B 62.3 45.4 4.8 43.6 0.6 32.2 Qwen2.5-VL-72B 98.8 81.9 42.9 80.6 67.8 50.0 GUI-specific Models (SFT) UI-TARS-2B 0.2 0.0 0.0 0.1 0.0 1.3 UI-TARS-7B 11.3 5.5 0.0 7.1 0.0 4.2 UI-TARS-72B 20.7 9.3 0.9 13.0 0.0 5.8 OS-Atlas-7B 37.3 27.7 2.4 26.1 0.0 23.0 ShowUI-2B 6.9 2.7 0.0 4.2 0.0 3.4 Aguvis-7B 8.3 4.5 0.0 5.3 0.0 29.7 GUI-specific Models (RL) GUI-R1-3B 55.4 37.2 3.0 37.8 0.0 24.3 GUI-R1-7B 65.4 45.0 9.3 46.3 0.0 28.0 InfiGUI-R1-3B 18.6 11.7 1.4 12.6 0.0 10.6 GUI-G1-3B 55.1 38.9 4.5 38.4 0.0 19.1 UI-TARS-1.5-7B 44.9 28.4 1.9 30.1 0.0 26.5 UI-Venus-Navi-7B 0.7 0.0 0.0 0.3 0.0 14.3 Ours RISK-SFT-7B 99.1 83.2 52.5 83.5 75.3 61.5 RISK-R1-7B 98.8 90.1 65.5 88.3 82.8 62.3
Offline Domain Evaluation. We compare RISK-R1 with commercial models, general open-source models, and GUI-specific models on RISK-Bench and OS-Genesis, as shown in Table 1. In single-step tasks on RISK-Bench, RISK-R1-7B outperforms GPT-4o by 6.8% and Qwen2.5-VL-72B by 7.7%. After RFT, RISK-R1-7B has a substantial increase of 6.9% and 22.6% in moderate and difficult tasks, respectively. This change reveals that level reweighting effectively guides the model to focus on challenging samples, enhancing its problem-solving capabilities. In multi-step tasks, RISK-R1-7B exceeds GPT-4o by 8.8%, indicating that multi-step trajectories in RISK-Data are beneficial for improving the task-level process understanding, while process reweighting emphasizing the importance of later steps in the trajectory also helps.
@@
Table 2: Performance comparison (success rate) of different models on RISK-Bench and OS-Genesis. The best and the second best results are highlighted in bold and underline, respectively. Our RISK-R1-7B surpasses all baselines across single-step and multi-step tasks.
Task GRPO DAPO GSPO Ours Single-step 84.9 85.7 86.3 88.3 Multi-step 75.4 77.6 79.4 82.8
Training Regimen Comparison. We benchmark our method against GRPO and its advanced variants, such as DAPO Yu et al. (2025) and GSPO Zheng et al. (2025), using the same training set (RISK-data) and the same SFT checkpoints (RISK-SFT-7B), summarized in Table 2. The results consistently demonstrate that our training regimen outperforms these strong baselines, confirming its effectiveness and specialized capability in risk management domains.
Table 3: Performance comparison of models with online evaluation. Compared with SOTA baselines, RISK-R1-7B achieves the highest task success rate while maintaining a competitive task completion rate.
Model Completion Rate Success Rate Commercial Models Qwen-VL-Max 85.2 66.2 General Open-source Models InternVL3-8B 0.0 0.0 Qwen2.5-VL-7B 8.3 46.4 Qwen2.5-VL-72B 88.7 68.9 GUI-specific Models (SFT) UI-TARS-7B 0.0 0.0 UI-TARS-72B 0.0 0.0 OS-Atlas-7B 4.4 37.2 ShowUI-2B 0.0 0.0 GUI-specific Models (RL) GUI-R1-7B 0.0 0.0 InfiGUI-R1-3B 0.0 0.0 GUI-G1-3B 0.0 0.0 UI-TARS-1.5-7B 0.0 0.0 Ours RISK-SFT-7B 86.1 67.0 RISK-R1-7B 87.6 70.5
Online Evaluation. For real-time multi-step decision-making evaluation, we use the Browser-Use framework to build a webpage interaction environment and compare RISK-R1 with various baselines. Different from offline evaluations, the model may encounter unseen webpages or changed page structures (even if the objective website is the same) during online evaluations, which poses a greater challenge to the model’s generalization ability and robustness. In Table 3, although the task completion rate of RISK-R1-7B is slightly lower than that of Qwen2.5-VL-72B, it achieves the highest task success rate of 70.5%, outperforming Qwen2.5-VL-72B by 1.6% and Qwen-VL-Max by 4.3%. This shows that RISK-R1 can effectively finish tasks in complex and dynamic real-world scenarios. More details on the cost and time analysis, as well as the error analysis, are provided in Appx. I and Appx. J.
@@ 5.3 Reward Design Analysis
Reward Effectiveness. We ablate each reward component on RISK-Bench (Table 4). The updated format reward $R_{for}$ improves single-step and multi-step accuracy by 1.0% and 2.7% by enforcing syntactic correctness and better task understanding. Process reweighting $\theta$ adds 0.6% and 1.2% by emphasizing critical steps in the trajectory. Level reweighting $w_{level}$ further brings +0.7% and +1.5% enhancements by focusing training on harder samples (No. 4 vs. No. 2). Stepwise accuracy reward $R_{step\_acc}$ adds +0.6% and +2.2% via finer-grained feedback. Overall, combining all components yields the best overall improvements, validating the effectiveness of our reward design.
@@
Fine-grained Feedback for Exploration. Stepwise accuracy reward is used in the early training stage to provide more fine-grained feedback for model to facilitate exploration. We analyze the impact of different reward settings, as shown in Figure 5. Leveraging stepwise accuracy reward in the early stage provides a faster reward enhancement by encouraging the model to learn from partially correct tool calls. Nevertheless, using stepwise reward throughout the entire training process does not yield better results than the solely binary accuracy reward, as the model may develop inertia under partial rewards. The optimal approach is to combine both reward types, using stepwise accuracy reward in the early stage and binary accuracy reward in the later stage, which effectively balances exploration and exploitation.
Table 4: Ablation study on reward design components in RISK-R1. Each component contributes positively to model performance.
No. $R_{for}$ $\theta$ $w_{level}$ $R_{step\_{acc}}$ Single-step Multi-step 1 84.9 75.4 2 85.9 (+1.0) 78.1 (+2.7) 3 86.5 (+1.6) 79.3 (+3.9) 4 86.6 (+1.7) 79.6 (+4.2) 5 87.7 (+2.8) 80.6 (+5.2) Ours 88.3 (+3.4) 82.8 (+7.4)
Figure 5: Stepwise accuracy reward analysis (Curves are smoothed). Peak performance is achieved at the combination of stepwise and binary accuracy rewards.
Table 5: Scaling law analysis of RISK on Qwen2.5-VL. RISK-R1 consistently improves the performance of Qwen2.5-VL across different model sizes.
Model Parameter Size Single-step Multi-step Qwen2.5-VL 3B 14.9 0.0 +RISK 3B 70.2 (+55.3) 62.5 (+62.5) Qwen2.5-VL 7B 43.6 0.6 +RISK 7B 88.3 (+44.7) 82.8 (+82.2) Qwen2.5-VL 32B 64.3 22.2 +RISK 32B 89.4 (+25.1) 84.9 (+62.7)
5.4 Scaling Law Analysis
To evaluate the scalability of RISK, we apply the RISK-R1 to Qwen2.5-VL models of different sizes and compare their performance on RISK-Bench, as shown in Table 5. RISK-R1 consistently enhances the performance of Qwen2.5-VL across all model sizes. Notably, the 3B model achieves the largest gains (+55.3% single-step, +62.5% multi-step), while the 32B model also improves substantially (+25.1% single-step, +62.7% multi-step). These findings show that RISK-R1 is an effective and scalable approach for enhancing GUI agent capabilities in e-commerce risk management tasks.